OpenStudy (anonymous):

Part 1: Using complete sentences, compare the key features and graphs of sine and cosine. What are their similarities and differences? Part 2: Using these similarities and differences, how would you transform f(x) = 2 sin(2x - π) + 3 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (camper4834):

similarities? their range is the same their period is the same they both repeat forever

OpenStudy (camper4834):

differences? the sine graph looks like a shifted version of the cosine graph

OpenStudy (camper4834):

im not sure what your instructor is looking for EXACTLY. Im sure you would have learned what words to use in class

OpenStudy (camper4834):

as for part two i'll get working on some graphs to show you

OpenStudy (camper4834):

but first i'll tell you this Acos(Bx-C)+D A will change the amplitude B will change the length of the period C will change where the period starts D will change the position of the cosine on the y axis

OpenStudy (camper4834):

since cosine is just a shifted version of sine we want A, B, and D to be exactly the same

OpenStudy (camper4834):

so lets start with 2cos(2x)+3

OpenStudy (camper4834):

the two graphs look almost right

OpenStudy (camper4834):

to shift the (red) cosine graph to look like the (blue) sine graph we have to change the C

OpenStudy (camper4834):

since i know the difference between sine and cosine will always be some fraction of pi or multiple of pi i can make an educated guess

OpenStudy (camper4834):

the distance LOOKS like 1 unit across so i can guess that the REAL distance is pi/4 since that equals around 1

OpenStudy (camper4834):

when shifting these kinds of graphs DISTANCE SHIFTED = -C/2 in our case distance shifted = pi/4 pi/4 = -C/2 pi/2 = -C -pi/2 = C

OpenStudy (camper4834):

so C equals -pi/2 plug that back into our equation we get 2cos(2x-(-pi/2))+3 which equals 2cos(2x+pi/2)+3

OpenStudy (camper4834):

since your question asks for there to be a negative sign there we can just do a little trick \[Acos(Bx \color{red}-C)+D\]

OpenStudy (camper4834):

adding or subtracting 2pi to C does nothing to the graph so if i just subtract 2pi \[\frac{ \pi }{ 2 }-2\pi=-\frac{ 3\pi }{ 2 }\] our new C is -3pi/2

OpenStudy (camper4834):

FINAL ANSWER TO PART 2 \[2\cos(2x-\frac{ 3\pi }{ 2 })+3\]